solution of 4s^3+12s^2+12s+4=0
Answer by richwmiller(17219)

4s^3+12s^2+12s+4=0
factor out 4
4*(s^3+3s^2+3s+1)=0
from this we see that the zeros are factors of 1
4*(s+1)^3=0
s=-1